I am just learning to do functions, domain.
When f(x)=x+1 and g(x)=2x find:
a) f(g(-3)=
so I did f=x+1
f= 2x+1
But I lose it for f(g(-3)
Is it f=x+1
f= (2x+1)-3
f= -6x+-3
I know this is wrong because I have the correct answer, but do not understand. The correct answer is 5,
please tell me what I m doing wrong?
Thank You
Marie

Answer by stanbon(75887)

When f(x)=x+1 and g(x)=2x find:
a) f(g(-3)=
----
Work these from the right to the left.
---
g(-3) = 2*-3 = -6
---
So, f(g(-3)) = f(-6) = -6+1 = -5
=====================================
cheers,
Stan H.